If 3.5 moles Ca(OH)2 are added to 380ml 1.3M NH4Cl how many moles of Ca(OH)2 will be left over?

NH4Cl + Ca(OH)2 --> NH3 + H2O + CaCl2

1
Expert's answer
2021-01-27T04:41:41-0500

Moles of NH"_4"Cl =Molarity*volume

moles=0.38*1.3=0.494

Here NH"_4" Cl is limiting reagent because it have less number of moles than calcium hydroxide.

1 mole calcium hydroxide react with 2 mole of NH"_4" Cl so,

remaining mole of calcium hydroxide =3.5-0.247=3.253 mole